a man is walking with a speed of 3m/s .raindrops are falling on the ground with speed of 3m/s what is the velocity of raindrops with respect to man?
 

Grade:9

3 Answers

Khimraj
5 years ago
angle between velocity of both is 90
so relative velocity will be 5\hat{i} + 5\hat{j} ( assuming downward as negative y axis)
so relative velocity = 5\sqrt{2} m/s
Hope it clears.
Sapna
5 years ago
Let velocity of man =v=3m/s
Let velocity of rain =V=3m/s 
resultant velocity is the velocity of rain wrt man=3^2+3^2=x^2 
=x=18^1/2
=4.24
Khimraj
5 years ago
sorry I take velocity  5m/s instead of 3m/s
so correct answer is 3\sqrt{2} m/s.